# PATCH: Learnable Tile-Level Hybrid Sparsity for LLMs

### v1

@@ 1 Introduction
To address the challenges of LLM pruning, while providing accelerated inference, we propose Pruning with a Learnable Tile-level Configuration for Hybrid Sparsity (PATCH). PATCH learns a hybrid mask that partitions each weight matrix into hardware-friendly tiles, designating each tile as either dense (0% sparsity) or 2:4 sparse (50% sparsity). This adaptive mask allows the matrix to realize an effective global sparsity ratio anywhere between 0% and 50%, balancing accuracy in critical regions with hardware-friendly sparsity elsewhere. This design unites the hardware acceleration benefits of 2:4 sparsity with the flexibility of unstructured allocation, allowing sparsity to adapt to the varying importance of different layers. By jointly optimizing the sparsity within 2:4 tiles and the tile-level patterns during training, PATCH achieves higher accuracy than uniform sparsity across layers. Moreover, for resource-constrained settings, we offer a variant of PATCH that tunes only the dense tiles while freezing the initial 2:4 mask fixed. Importantly, PATCH is compatible with tile-level sparsity acceleration libraries and compilers such as STOICC (Rafii et al., 2025), making it the first hybrid sparsity method to demonstrate practical speedups. For example, on LLaMA-2 7B running on a consumer-grade A6000 GPU, PATCH achieves 1.18–1.38 end-to-end speedup over the dense baseline while improving accuracy by 0.37%–2.96% compared to the state-of-the-art 2:4 pruning method, MaskLLM.
@@ 5.2 Understanding the components of PATCH
Global | Layer-wise | ||
| 45% | 14.57 | 15.17 | ||
| 35% | 13.84 | 14.48 | ||
@@
| Sparsity (0.5B) | MaskLLM |
Wanda | Magnitude | PATCH | ||
| 45% | 15.06 | 21.84 | 21.83 | 21.33 | 14.57 | ||
| 35% | 14.55 | 17.29 | 17.96 | 19.90 | 13.84 | ||
@@ 5.3 Combination with other compression methods
Table 6 reports results on LLaMA-2 7B and LLaMA-3.1 8B, showing that combining sparsity, quantization, and low-rank approximation enables controllable tradeoffs between compression ratio and model quality, highlighting PATCH ’s versatility.
| Sparsity | Method | Pattern | LoRA | LLaMA-2-7B | LLaMA-3.1-8B | ||
| Acc (% ) | PPL () | Acc (% ) | PPL () | ||||
| 0% | Dense | - | - | 56.44 | 5.12 | 59.85 | 5.84 |
| 50% | MaskLLM | 2:4 | - | 47.98 | 7.64 | 51.12 | 9.92 |
@@ Appendix D Implementation details and hyperparameters
Training is accelerated via data parallelism across a single node with 4 H100 GPUs. In this setup, PATCH requires 4.5 and 6 GPU hours on the 0.5B and 1B models, respectively, while PATCH requires 21 and 24 GPU hours on the 7B and 8B models.
@@ Appendix G Comparison with unstructured sparsity
Appendix H Language model usage in paper
@@
Appendix I Reproducibility statement

### v3

@@ 1 Introduction
To address the challenges of LLM pruning, while providing accelerated inference, we propose Pruning with a Learnable Tile-level Configuration for Hybrid Sparsity (PATCH). PATCH learns a hybrid mask that partitions each weight matrix into hardware-friendly tiles, designating each tile as either dense (0% sparsity) or 2:4 sparse (50% sparsity). This adaptive mask allows the matrix to realize an effective global sparsity ratio anywhere between 0% and 50%, balancing accuracy in critical regions with hardware-friendly sparsity elsewhere. This design unites the hardware acceleration benefits of 2:4 sparsity with the flexibility of unstructured allocation, allowing sparsity to adapt to the varying importance of different layers. By jointly optimizing the sparsity within 2:4 tiles and the tile-level patterns during training, PATCH achieves higher accuracy than uniform sparsity across layers. Moreover, for resource-constrained settings, we offer a variant of PATCH that tunes only the dense tiles while freezing the initial 2:4 mask. Importantly, PATCH is compatible with tile-level sparsity acceleration libraries and compilers such as STOICC (Rafii et al., 2025), making it the first hybrid sparsity method to demonstrate practical speedups. For example, on LLaMA-2 7B running on a consumer-grade A6000 GPU, PATCH achieves 1.18–1.38 end-to-end speedup over the dense baseline while improving accuracy by 0.37%–2.96% compared to the state-of-the-art 2:4 pruning method, MaskLLM.
@@ 5.2 Understanding the components of PATCH
| Sparsity (0.5B) | Global | Layer-wise |
| 45% | 14.57 | 15.17 |
| 35% | 13.84 | 14.48 |
@@
| Sparsity (0.5B) | MaskLLM | SparseGPT (w/o weight update) | Wanda | Magnitude | PATCH |
| 45% | 15.06 | 21.84 | 21.83 | 21.33 | 14.57 |
| 35% | 14.55 | 17.29 | 17.96 | 19.90 | 13.84 |
@@ 5.3 Combination with other compression methods
Table 6 reports results on LLaMA-2 7B and LLaMA-3.1 8B, comparing PATCH and MaskLLM under 4-bit weight-only quantization, as well as an additional setting that combines our method with a low-rank adapter (of 10% of the weight’s rank). These results show that sparsity, quantization, and low-rank approximation can be composed to achieve controllable tradeoffs between compression and model quality, and that our approach integrates seamlessly with both techniques within broader compression pipelines.
| Sparsity | Method | Bit | LoRA | LLaMA-2-7B | LLaMA-3.1-8B | Comp. | ||
| Acc (% ) | PPL () | Acc (% ) | PPL () | Ratio | ||||
| 0% | Dense | - | - | 54.61 | 5.12 | 60.31 | 5.84 | 1x |
| 50% | MaskLLM | 4 | - | 47.98 | 7.64 | 51.12 | 9.92 | 5.33x |
@@ Appendix D Implementation details and hyperparameters
Training is accelerated via data parallelism across a single node with 4 H100 GPUs. In this setup, PATCH requires 18 and 24 GPU hours on the 0.5B and 1B models, respectively, while PATCH requires 84 and 96 GPU hours on the 7B and 8B models.
@@ Appendix G Comparison with unstructured sparsity
Appendix H Layer-wise sparsity distribution comparison with other work
We compare the sparsity allocation learned by PATCH with OWL (Yin et al., 2025) and AlphaPruning (Lu et al., 2024). To ensure a robust baseline, we performed extensive hyperparameter sweeps for both methods and selected the configurations that achieved the best perplexity when applying Wanda as the underlying pruning operator:
OWL: We swept and .
AlphaPruning: We swept 20 values of the temperature parameter between 0 and 0.5.
A critical distinction is that OWL and AlphaPruning are primarily designed for unstructured sparsity. While OWL includes an N:M structured variant, it is restricted to a fixed 50% global sparsity. Furthermore, OWL allocates sparsity at the block level (assigning uniform sparsity to all matrices within a Transformer block). Although the authors propose a weight-wise variant, they report, and our experiments confirm, that it yields inferior performance. In contrast, PATCH operates with tile-level granularity within a semi-structured constraint, offering a unique combination of hardware acceleration and fine-grained control.
As seen in Figures 6 and 7, both baselines produce relatively flat sparsity distributions across the model. AlphaPruning exhibits only minor fluctuations in the middle layers, whereas PATCH discovers distinct, highly non-uniform patterns (e.g., preserving Attention layers while aggressively pruning MLP blocks).
For OWL, we performed a sweep over the parameter in and the parameter in for each model. For AlphaPruning, we swept 20 values of . For both methods, we selected the hyperparameters that achieved the best perplexity using Wanda as the pruning metric.
Appendix I Variation across seeds
We evaluate PATCH on Qwen-2.5 0.5B and Llama-3.2 1B across different seeds. As shown in Table 17, model performance remains consistent across seeds.
| Model | Sparsity (%) | Seed 0 (default) | Seed 25 | Seed 26 | Seed 42 |
| Qwen-2.5 0.5B | 25 | 13.47 | 13.41 | 13.38 | 13.36 |
| 35 | 13.84 | 13.89 | 13.85 | 13.84 | |
| 45 | 14.56 | 14.59 | 14.61 | 14.49 | |
| Llama-3.2 1B | 25 | 11.00 | 11.09 | 11.03 | 11.21 |
| 35 | 11.67 | 11.72 | 11.56 | 11.86 | |
| 45 | 12.23 | 12.32 | 12.26 | 12.55 |
In addition, we include the corresponding sparsity allocations across layers (Figure 9) and across individual weight matrices for both Llama-3.2 1B (Figure 10) and Qwen-2.5 0.5B (Figure 11). From these, we observe the following:
At the global block level, the optimization consistently identifies the middle Transformer blocks as the most redundant (receiving the highest sparsity), while the initial and (especially) final blocks are pruned less.
At the weight-matrix level, The allocation of sparsity between attention and MLP modules remains stable. For example, in Llama-3.2 1B (Figure 10), the Key and Value matrices consistently remain dense across all seeds, while the Query and MLP matrices absorb the majority of the sparsity
While the specific tile indices may vary slightly due to the stochastic sampling, the macroscopic pruning strategy learned by PATCH is highly reproducible and robust to initialization.
Appendix J Fine-Tuning after mask training
In this section, we present results from fine-tuning the remaining unpruned weights after mask training. We conducted a brief 5.6M-token run on the SlimPajama dataset so that the fine-tuning phase matches the mask search of PATCH.
Fine-tuning yields a consistent improvement in average zero-shot accuracy (e.g., +0.8% for Llama-3.2 1B at 25% sparsity and +0.6% for Qwen-2.5 0.5B at 45% sparsity), as seen on on Table 18. Interestingly, we observe a slight degradation in perplexity. We attribute this to the limited calibration data (5.6M tokens) compared to the trillions of tokens seen during pre-training; short fine-tuning can sometimes slightly drift the language modeling distribution while sharpening downstream task performance.
These results confirm that PATCH creates a high-quality sparsity topology that serves as a strong foundation. While the mask alone delivers state-of-the-art performance, subsequent fine-tuning, even with a limited budget, can further recover accuracy, offering a flexible path for users with additional compute resources.
| Model | Sparsity | Method | Wiki PPL () | Avg Acc (% ) |
| Qwen-2.5 0.5B | 45% | PATCH | 14.56 | 40.29 |
| PATCH + FT | 14.96 | 40.87 | ||
| 35% | PATCH | 13.84 | 41.15 | |
| PATCH + FT | 14.32 | 41.59 | ||
| 25% | PATCH | 13.47 | 42.39 | |
| PATCH + FT | 13.85 | 42.55 | ||
| LLaMA-3.2 1B | 35% | PATCH | 11.67 | 42.72 |
| PATCH + FT | 12.02 | 43.50 | ||
| 25% | PATCH | 11.00 | 43.81 | |
| PATCH + FT | 11.36 | 44.61 |
Appendix K Impact of sparse fine-tuning under fixed compute budget
To assess whether the performance gains of PATCH are solely due to training, we conducted a controlled experiment comparing PATCH against fine-tuned (FT) one-shot baselines under a strictly fixed compute budget. We fine-tuned the weights of Wanda and SparseGPT (2:4 sparsity) models on the SlimPajama dataset for 5.6M tokens. We trained the PATCH masks for the equivalent compute budget (matching the 5.6M token run).
As shown in Table 19, PATCH consistently outperforms the fine-tuned baselines, even when the baselines are allowed to update their weights. For example, on LLaMA-3.2 1B, PATCH at 45% sparsity achieves 42.08% accuracy, surpassing SparseGPT + FT (41.74%) and Wanda + FT (41.27%). This demonstrates that under a fixed compute budget, learning a flexible, non-uniform sparsity mask yields better performance than fine-tuning weights with a rigid, uniform mask.
| Model | Method | Sparsity (%) | PPL () | Avg Acc () |
| Qwen-2.5 0.5B | Wanda (2:4) | 50 | 72.48 | 32.97 |
| Wanda + FT | 50 | 15.21 | 39.98 | |
| SparseGPT (2:4) | 50 | 36.59 | 34.81 | |
| SparseGPT + FT | 50 | 16.98 | 38.89 | |
| PATCH | 45 | 14.56 | 40.29 | |
| PATCH | 35 | 13.84 | 41.15 | |
| PATCH | 25 | 13.47 | 42.39 | |
| LLaMA-3.2 1B | Wanda (2:4) | 50 | 78.18 | 31.61 |
| Wanda + FT | 50 | 13.98 | 41.27 | |
| SparseGPT (2:4) | 50 | 32.73 | 35.55 | |
| SparseGPT + FT | 50 | 13.54 | 41.74 | |
| PATCH | 45 | 12.23 | 42.08 | |
| PATCH | 35 | 11.67 | 42.72 | |
| PATCH | 25 | 11.00 | 43.81 |
Appendix L Language model usage in paper
@@ Appendix I Variation across seeds
Appendix M Reproducibility statement